OpenStudy (anonymous):

Can someone check my work? 1. Find the period for f(x) = -7sin5x Answer: 2pi/5?

Nnesha (nnesha):

\[\huge\rm y=a \sin(bx-c)\] where \[\left| a \right|\] = amplitude to find period use the formula \[\huge\rm \frac{ 2\pi }{ b }\]

OpenStudy (anonymous):

So then my answer is correct?

Nnesha (nnesha):

opps didn't see the answer yes looks correct

OpenStudy (anonymous):

Thank you!

Nnesha (nnesha):

yw \o^_^o/

OpenStudy (anonymous):

Are you really good with trigonometry? I may need some help with some more questions of this type.

Nnesha (nnesha):

not sure depends on the question i'm not good 't graphing sin,tancos functions sorry

OpenStudy (anonymous):

Aw, okay. Well I'll post another question and tag you in it. :)